Mention the commutativity, associative and distributive properties of rational numbers. Also, check a × b = b × a and a + b = b + a for a = ½ and b = ¾

Grade:12

3 Answers

Harshit Singh
3 years ago
Dear Student

Commutative property:
For any two rational numbers a and b, a + b = b + a.
For any two rational numbers a and b, a × b = b × a.
Associative Property:
For any three rational numbers a, b and c,
(a + b) + c = a + (b + c)
Distributive property states that for any three numbers x, y and z,
x × ( y + z ) = (x × y) + ( x × z)

a*b = b*a
a*b = ½ * ¾ = 3/8
b*a = ¾ * ½ = 3/8
a + b = ¾ + ½ = 5/4
b + a = ½ + ¾ = 5/4

Thanks
Yashraj
3 years ago
  1. Practice previous year questions many times..
  2. Appear for a test series and analyse it properly (Giving 2–3 hours)
  3. Revise your notes multiple times till they enter your soul :)
  4. Find your strong points and chapters, work on them..
  5. Solve other people’s doubts. It makes your own concepts very clear
Prawin Gupta
2 years ago
Commutative Proprety
In this property,we can change the order if the sign between them are of 'x' and '+' 
Like 6+8=8+6
        14=14
 
Associative Property
In this property we can shift the brackets from left to right or right to leftthe sign between them would also be of '+' or 'x'
Like (1x2)X3=1x(2x3)
        2x3.       =1x6
         6.           =6
 
Distributive Property
In this we can take a common number out of the bracket (only work if the sign is of 'x' or'+'
Like 3x2+3x5
         3(2+5)
          3(7)
            21